OpenStudy (anonymous):

Find the equation for the line that passes through the points (-1,-1) and (-5,5)

OpenStudy (imtiaz7):

hre we first find its slope by forumula m=y2-y1/x2-x1 by solving this we get the slope m=-3/2 : then

OpenStudy (anonymous):

Is that all?

OpenStudy (imtiaz7):

we'll use point slope form that is y-y1=m(x-x1) so by putting the values we get the equation: 3x+2y+5=0 thats the final equation!

OpenStudy (anonymous):

Ohh ok

OpenStudy (imtiaz7):

do you understand it?

OpenStudy (anonymous):

Yep thanks

OpenStudy (imtiaz7):

Hmm, you're forgetting something..

OpenStudy (anonymous):

Im new to this so idk

OpenStudy (imtiaz7):

:) its ok and lots of welcome here